Kara Ries, Tammy Bax, and Joe Thomas invested $42,000, $58,000, and $66,000, respectively, in a partnership. During its first calendar year, the firm earned $400,500.

Prepare the entry to close the firm’s Income Summary account as of its December 31 year-end and to allocate the $400,500 net income to the partners under each of the following separate assumptions:

(2) The partners agreed to share income and loss in the ratio of their beginning capital investments. (Do not round intermediate calculations. Round final answers to the nearest whole dollar.)
Answer #1
Debit Credit
Dec 31 Income summary 400500
         Kara Ries, Capital 101331 =400500/(42000+58000+66000)*42000
        Tammy Bax, Capital 139934 =400500/(42000+58000+66000)*58000
         Joe Thomas, Capital 159235 =400500/(42000+58000+66000)*66000